tony's age is 17 and his father's age is 50. How many years ago was the father four times as old as Tony?
Answer by addingup(3677)

50-x = 4(17-x)
50-x = 68-4x
3x = 18
x = 6 Six years ago Tony's father was 4 times Tony's age:
Tony today: 17
Tony 6 yrs ago: 17-6 = 11
11*4 = 44 Father is 50 and 50-6 = 44 Correct